# Complex Saddles of Charged-AdS Gravitational partition function

### v1

@@ Abstract
In this paper, we consider the Euclidean partition function of charged and uncharged AdS black hole geometries in $d+1$-dimensions. It is seen that the partition function can be reduced to a one-dimensional integral, which can be investigated using methods of Picard-Lefschetz. The saddles of the system correspond to either naked-singular geometry, thermal-AdS, small-, intermediate- or large-sized black hole for different ranges of parameter space. These are solutions of Einstein’s equation, which are dominant saddles in the partition function in various regimes of parameter space. A naive analysis of the partition function involving these saddles would lead to conflicts with the standard understanding of black hole thermodynamics and also with AdS/CFT. However, when the partition function is analysed using Picard-Lefschetz, it is seen that naked-singular geometries don’t contribute. The saddles corresponding to them are irrelevant, aligning well with the Cosmic Censorship hypothesis. Saddles corresponding to negative specific heat are either small- or intermediate-sized black holes. Although they are relevant in the partition function but are sub-dominant. They also drop out under homology averaging. Saddles corresponding to only non-negative specific heat contribute to the Euclidean partition function. Finally, we analyze the allowability of these complex geometries using the KSW criterion.
@@ I Introduction
where the trace $\rm Tr$} is defined as the summation over all the energy eigenstates. In the case of gravity, in a sense the same is true as pointed out long ago by Gibbons and Hawking Gibbons:1976ue , where the gravitational path-integral $Z(\beta)$ is defined as path-integral over some class of metrics with Euclidean signature boundary $S^{1}\times\mathbb{Y}$ where $S^{1}$ has proper length $\beta$. It is well-known that path-integrals get dominant contribution from saddle configurations, which are configurations satisfying equation of motion of the action describing the system. This thing works well in non-gravitational cases where the saddles correspond to configuration with least energy. In gravitational systems, things are not so straight forward.
@@
Euclidean gravitational path-integral suffers from conformal factor problem, where the path-integral over the conformal factor is unbounded from below Gibbons:1978ac . This issue arises even before one can address choice of integration contour. A simple treatment proposed earlier was to rotate the contour of integration of the scale-factor, which although is adhoc but does the task. This procedure has been utilized by several authors giving physically satisfying results Allen:1984bp ; Prestidge:1999uq ; Kol:2006ga ; Headrick:2006ti ; Monteiro:2008wr ; Monteiro:2009tc ; Marolf:2018ldl where the choice of contour is adhoc lacking justification from first principles.
@@
A better proposal is to deal with the Lorentzian path-integral where one doesn’t hit the conformal factor problem directly in the sense that the path-integral is defined over Lorentzian geometries with the contour of integration taken to be defined by real Lorentzian-signature metrics. The notion then is carefully scrutiny the possible deformation of integration contour in the complex plane to come up with an equivalent “Euclidean” path-integral which right from beginning comes along with a specific integration contour where meaningful computation can be performed. This approach has been considered by various authors in both continuum and discrete formulation of path-integral of gravity Schleich:1987fm ; Mazur:1989by ; Hartle:2020glw ; Giddings:1989ny ; Marolf:1996gb ; Dasgupta:2001ue ; Ambjorn:2002gr ; Feldbrugge:2017kzv ; Feldbrugge:2017fcc ; Ailiga:2023wzl ; Ailiga:2024wdx ; Ailiga:2025fny .
@@
The benefit of using a Lorentzian formulation of path-integral for gravity is that it is not ruled out for obvious reasons like un-boundedness or blowups happening in Euclidean path-integral of gravity. This is because the Lorentzian gravitational action $I_{\rm grav}$ for smooth real Lorentzian metrics is purely real, hence implying that integrand of the path-integral $\exp(iI_{\rm grav})$ is a pure phase leading to an oscillatory integral! Is there a way to define the Euclidean path-integral $Z(\beta,\cdots)$ (where the $(\cdots)$ represent charges, angular momentum, etc.) via Lorentzian path-integral by some transformation? This was investigated in Marolf:2022ybi , where the author proposes to consider Lorentzian path-integral as path-integral over all Lorentzian metrics including co-dimension-2 singularities, and uses a smearing function $f_{\beta}(T)$ to covert the Lorentzian path-integral to a “Euclidean” one. The smearing function is defined by
$\int_{-\infty}^{\infty}{\rm d}Tf_{\beta}(T)\,\,e^{-iT\omega}=e^{-\beta\omega}\,$, (2)
@@
for all $\omega$ above the ground state of Hamiltonian. This immediately gives that the Euclidean partition function can be written as
@@
where $S={\rm Area}(\gamma)/4G_{N}$ and $Z_{L}(T;S)$ is the path-integral over Lorentzian metrics of fixed $S$ (or fixed area of the co-dimension-2 surface $\gamma$). Saddle configurations of $Z_{L}(T;S)$ then becomes constrained-saddles of the $Z_{L}(T)$. These constrained saddles can become relevant depending on the integration contour and can contribute non-trivially in the path-integral once the integration contour is deformed using Picard-Lefschetz methodology Marolf:2022ybi ; Chen:2025leq .
@@
The essence of this procedure is that we donot deal with the conformal factor problem of the Euclidean path-integral, instead uses Lorentzian path-integral to define a Euclidean one via an integral transform, which for the case of black hole geometries reduces to known partition function.
Once we have comfortably defined the Euclidean partition function $Z(\beta)$ for the case of black hole geometries, we proceed to study it in detail for the case of AdS-Schwarzschild non-charged and charged black holes in arbitrary dimensions. Our starting point in this study is eq. (7) and eq. (8). This is a one/two-dimensional integral which can be evaluated by analysing it in complex plane. As this targets geometries which correspond to black hole solutions, it is worth exploring what are the possible saddles of the system and what do they correspond to. Furthermore, it is crucial to investigate which saddles give dominant contribution to the partition function for various values of parameter $\beta$ (and charge $Q$ in case of charged BHs) and whether they are compatible with the known results of AdS/CFT for various dimensions.
It may happen that certain saddles which are dominant in the path-integral for a range of external parameters $\beta$ and charge $q$ (or chemical potential $\Phi$), need not be compatible with the known wisdom of AdS/CFT. In such cases it is worth asking whether such saddles become “relevant” in the partition function when the integration mentioned in eq. (7) is performed in the complex plane using Picard-Lefschtz methods (Morse theory). An interesting situation seen in the recent work Mahajan:2025bzo ; Singhi:2025rfy that there are “relevant” saddles not contributing in the partition function. Contributions from such saddles are noticed to drop out in the partition function due to homology averaging. This paper investigates such situations in more detail for the case of charged and un-charged black holes in $d+1$-dimensions.
The outline of the paper is as follows: In Sec. II, we revisit the puzzle and findings reported in Mahajan:2025bzo for asymptotically AdS Schwarzschild spacetimes in arbitrary dimensions. Here, we discuss the implementation of Picard Lefschetz and homology averaging for generic complex saddles, along with associated thermodynamic stability analysis. In Sec. III, we extend this study to the charged AdS case, considering arbitrary dimensions for the grand canonical ensemble (fixed potential) and some specific dimensions for the canonical ensemble (fixed charge). Sec. IV is devoted for the analysis of weaker verison of KSW criterion and its compatibility with the puzzle raised in the earlier sections. Finally, summary of our findings and conclusions are presented in section V.
II Schwarzschild AdS black hole in $d+1$ dimension
@@
This is solution to equation of motion following from Einstein-Hilbert action with appropriate surface terms and counterterms to cancel the large volume divergence Skenderis:2002wp . For $\mu=0$, it describes a thermal AdS solution without any horizon. However, when $\mu>0$, we get a horizon at $r_{+}$, which is the largest positive real root of $f(r_{+})=0$ and $_{+}\leq r<\infty. rr$$\geq r_{+}$ ensures the existence of a timelike killing vector everywhere outside the horizon, regularity at the horizon enforces $\beta$ to be $4\pi/f^{\prime}(r_{+})$. The entropy $S$ and Energy $E$ as function of $r_{+}$ are given by
$S(r_{+})=\frac{\omega r_{+}^{d-1}}{4G},\hskip 11.38109ptE(r_{+})=\frac{\omega(d-1)}{16\pi G}\left(r_{+}^{d-2}+\frac{r_{+}^{d}}{l^{2}}\right).$ (10)
@@
$\beta_{\rm max}=\frac{2l\pi}{\sqrt{d(d-2)}},$ (14)
@@
$\begin{split}K(r)=\frac{d(d-1)^{2}(d-2)\mu^{2}}{r^{2d}}+\frac{2d(d+1)}{l^{4}}\,,\quad{\rm where}\quad\mu^{2}=r_{+}^{2(d-2)}\bigl(1+\frac{r_{+}^{2}}{l^{2}}\bigr)^{2}\,.\end{split}$ (17)
From eq (17), it is evident that at $r=0$, $K(r=0)$ blows up in $d\geq 3$, signifying the presence of curvature singularity irrespective of whether $r_{+}$ is real or complex (which happens for $\beta>\beta_{\rm max}$). For $\beta>\beta_{\rm max}$, these saddle configurations are singularities without a horizon, a naked singularity. Such naked singularities are solutions of Einstein’s equations.
In the following subsections, we will study the partition function for various regimes of $\beta$, analysing the saddle configurations, their dominance and relevance, and compute the partition function using the methods of Picard-Lefschetz. To do so, we will need to compute study the argument of the exponent in eq. (11) at the various saddles. $\mathcal{I}(r_{+})$ at various saddles is given by
[…]
@@ II.1 $Z(\beta)$ for $\beta>\beta_{\rm max}$
In this regime, we have one real saddle and two complex saddle configurations. The real saddle corresponds to thermal AdS geometry, while the complex-conjugate saddle correspond to naked singular geometries. As noticed in Mahajan:2025bzo , these complex saddles give dominant contribution in the partition function for certain range of values of $\beta$, as can be seen by computing $\mathcal{I}(r_{+})$. Naively, it will be expected that being dominant saddles, they are the ones whose contribution will be most important in the classical limit. However, the situation changes if such saddles, despite being dominant becomes irrelevant in the computation of the partition function using Picard-Lefschetz methods. This situation was analysed for the case of $d=4$ in Mahajan:2025bzo . In our paper we examine the situation in arbitrary dimensions.
Figure 1: Steepest descent/ascent lines are plotted in black/red for $AdS_{6}$, for $\beta=5\beta_{\rm max}$, $l=1$. Thermal-AdS is at $r_{+}=0$ and $r_{+}^{\pm}$ are naked singularities. The purple arrow is the direction of integration cycle. Steepest ascent curves from these complex saddles don’t intersect the real line and hence don’t contribute to the partition function.
In order to compute the partition function for $\beta>\beta_{\rm max}$, we start by analysing $Re(\mathcal{I}(r_{+}))$ at various saddles. It will indicate the saddles which will become dominant as $\beta$ varies. It is seen from eq. (18) that for arbitrary $d$, $\mathcal{I}(r_{+})$ for thermal AdS is always zero. For $\beta>\beta_{\rm max}$, the other two saddles $r_{+}^{\pm}$ are complex conjugate and correspond to naked singularities. For them $\mathcal{I}(r_{+})$ is complex as seen from eq. (18). To analyse the dominance, one needs to extract the real part of $Re[\mathcal{I}(r_{+})]$ for each of these complex saddles. If the $Re[\mathcal{I}(r_{+})]>0$ for certain values of $\beta$, then these naked singularities become dominant over the thermal-AdS saddle, which will be in conflict with the results of AdS/CFT. For $\beta>\beta_{\rm max}$, the $Re[\mathcal{I}(r_{+}^{\pm})]$ is given by
@@
It is easy to see from here that for certain values of $\beta$, say for example large-$\beta$ ($\beta\to\infty$), $\cos((d-1)\theta)=0$ or $\pm 1$, while $\sin((d-1)\theta)=\pm 1$ or 0 (depending on the dimension $d$), respectively. In the case when it is 1-1 (for example, $d=4$), then naked singularities are dominant in the partition function! More generically, for large-$\beta$, when $d=4n,n\in(1,2,3,..)$, $Re[\mathcal{I}(r_{+}^{\pm})]\sim+(\beta/\beta_{\rm max})>0$ thereby claiming dominance over thermal-AdS. This was already noticed in Mahajan:2025bzo for $AdS_{5}(n=1)$. When d=4n1d=4n-1, for large-$\beta$, $Re[\mathcal{I}(r_{+}^{\pm})]<0$ (independent of $\beta$) and thermal-AdS remains a dominant saddle. However, for $d=4n+1$, $Re[\mathcal{I}(r_{+}^{\pm})]>0$ and is independent of $\beta$, thereby again dominating over the thermal-AdS saddle. For $d=4n+2$, $Re[\mathcal{I}(r_{+}^{\pm})]\sim-(\beta/\beta_{\rm max})$ for large-$\beta$ thereby making thermal-AdS dominant saddle. We summarize this in the table 1.
@@
Table 1: Large $\beta$ ($\gg\beta_{\rm max}$) behaviour of the partition function at arbitrary $d$. $n\in(1,2,3,\cdots)$ at various saddles showing the dominance of various saddles.
On doing a numerical plot of $Re[\mathcal{I}(r_{+}^{\pm})]$ as a function of arbitrary $\beta>\beta_{\rm max}$, we notice from fig. (2) the values of $\beta$ for which naked singularities give dominant contribution in the partition function.
Figure 2: Real part of $\mathcal{I}$ as a function of $\beta$ ($\beta_{\rm max}$) at various saddles for $d=4$ and $d=7$. $Re[\mathcal{I}(r_{+}^{\pm})]>0$ for complex saddles (naked-singular geometries) imply that for these values of $\beta$ naked singularities are dominant over the Thermal-AdS saddle.
From this discussion, it is clear that for certain values of $\beta$ for the case of $\beta>\beta_{\rm max}$, naked singular geometries, which are complex saddles, are dominant. Naively, one would expect that they would control the nature of the partition function in the classical limit, as they are most dominant, though they correspond to complex metrics ($r_{+}$ being complex implies $\mu$ is complex). Occurrences of complex metrics are not new in the literature. They have shown up in studies of quantum cosmology and black holes Feldbrugge:2017kzv ; Maldacena:2024uhs ; Ivo:2024ill ; Ailiga:2024mmt ; Ailiga:2023wzl ; DiTucci:2020weq ; Lehners:2021mah ; Jonas:2022uqb ; Nishimura:2024one ; Hertog:2015nia ; Cabo-Bizet:2020ewf and real-time holography as well Skenderis:2008dh ; Glorioso:2018mmw ; Jana:2020vyx , where they have lead to physically meaningful results. However, in our case if they are to be considered, then our findings will be in tension with the understanding of black hole thermodynamics. Furthermore, if these additional saddles are considered, new phase transitions will occur among these complex saddles. This will also be in conflict with AdS-CFT, where it is well understood that thermal AdS reproduces results consistent with the dual CFT in the low temperature regime Witten:1998zw ; Witten:1998qj ; Aharony:2003sx .
This is the puzzle first raised in the note Mahajan:2025bzo , where naive computation of the partition function based on saddle dominance lead to tension with AdS/CFT. However, if the partition function is computed carefully, making use of Picard-Lefschetz methods and tools from Morse theory, then the puzzle is resolved on its own, as was noticed in Mahajan:2025bzo . In this paper, we study this situation for generic dimensions and analytically show that complex saddles corresponding to naked-singular geometries become irrelevant in the path-integral and hence don’t contribute.

### v2

@@ Abstract
In this paper, we consider the Euclidean partition function of uncharged and charged $AdS_{d+1}$ black hole geometries in canonical and grand canonical ensemble for $d\geq 3$. It is seen that the partition function can be reduced to a one-dimensional integral, which can be investigated using methods of Picard-Lefschetz. The saddles of the system correspond to either naked-singular geometry, thermal-AdS, small-, intermediate- or large-sized black hole for different ranges of parameter space. These are solutions of Einstein’s equation, which are dominant saddles of the partition function in various regimes of parameter space. A naive analysis of the partition function involving these saddles would lead to conflicts with the standard understanding of black hole thermodynamics and also with AdS/CFT. However, when the partition function is analysed using Picard-Lefschetz, it is seen that naked-singular geometries turn out to be irrelevant and therefore do not contribute. This also aligns well with the Cosmic Censorship hypothesis. Depending on the ensemble, saddles corresponding to negative specific heat are either small- or intermediate-sized black holes. Although they are relevant in the partition function but are sub-dominant. They drop out under homology averaging. Only saddles corresponding to non-negative specific heat contribute to the Euclidean partition function. Finally, we analyze the allowability of these complex geometries using the KSW criterion.
@@ I Introduction
where the trace $\rm Tr$} is defined as the summation over all the energy eigenstates. In the case of gravity, in a sense the same is true as pointed out long ago by Gibbons and Hawking Gibbons:1976ue , where the gravitational path-integral $Z(\beta)$ is defined as path-integral over some class of metrics with Euclidean signature boundary $S^{1}\times\mathbb{Y}$ where $S^{1}$ has proper length $\beta$. It is well-known that path-integrals get dominant contribution from saddle configurations, which are the configurations satisfying the onshell equation of the action describing the system. This thing works well in non-gravitational cases where the saddles correspond to configuration with least energy. In gravitational systems, things are not so straightforward.
@@
Euclidean gravitational partition function suffers from the conformal factor problem, where the path-integral over the conformal factor is unbounded from below Gibbons:1978ac . A simple treatment proposed earlier was to rotate the contour of integration of the conformal-factor, which although is adhoc but does the task. This procedure has been utilized by several authors giving physically satisfying results Allen:1984bp ; Prestidge:1999uq ; Kol:2006ga ; Headrick:2006ti ; Monteiro:2008wr ; Monteiro:2009tc ; Marolf:2018ldl where the choice of contour is ad hoc lacking justification from first principles.
@@
A better proposal is to deal with the Lorentzian path-integral where one doesn’t encounter the conformal factor problem directly in the sense that the path-integral is defined over Lorentzian geometries with the contour of integration taken to be defined by real Lorentzian-signature metrics. The notion is then carefully scrutinize the possible deformation of integration contour in the complex plane to come up with an equivalent “Euclidean” path-integral which right from beginning comes along with a specific integration contour where meaningful computation can be performed. This approach has been considered by various authors in both continuum and discrete formulation of path-integral of gravity Schleich:1987fm ; Mazur:1989by ; Hartle:2020glw ; Giddings:1989ny ; Marolf:1996gb ; Dasgupta:2001ue ; Ambjorn:2002gr ; Feldbrugge:2017kzv ; Feldbrugge:2017fcc ; Ailiga:2023wzl ; Ailiga:2024wdx ; Ailiga:2025fny .
@@
The benefit of using a Lorentzian formulation of path-integral for gravity is that it is not ruled out for obvious reason like unboundedness happening in Euclidean gravity. This is because the Lorentzian gravitational action $I_{\rm grav}$ for smooth real Lorentzian metrics is purely real, hence implying that integrand of the path-integral $\exp(iI_{\rm grav})$ is a pure phase leading to an oscillatory integral! Is there a way to define the Euclidean path-integral $Z(\beta,\cdots)$ (where the $(\cdots)$ represent charges, angular momentum, etc.) via Lorentzian path-integral by some transformation? This was investigated in Marolf:2022ybi , where the author proposes to consider Lorentzian path-integral as path-integral over all Lorentzian metrics including co-dimension-2 singularities, and uses a smearing function $f_{\beta}(T)$ to convert the Lorentzian path-integral to a “Euclidean” one. The smearing function is defined by
$\int_{-\infty}^{\infty}{\rm d}Tf_{\beta}(T)\,\,e^{-iT\bar{\omega}}=e^{-\beta\bar{\omega}}\,$, (2)
@@
for all $\bar{\omega}$ above the ground state of Hamiltonian. This immediately gives that the Euclidean partition function can be written as
@@
where $S={\rm Area}(\gamma)/4G_{N}$ and $Z_{L}(T;S)$ is the path-integral over Lorentzian metrics of fixed $S$ (or fixed area of the co-dimension-2 surface $\gamma$). Saddle configurations of $Z_{L}(T;S)$ then become constrained saddles of the $Z_{L}(T)$. These constrained saddles can become relevant depending on the integration contour and can contribute non-trivially to the path integral once the integration contour is deformed using the Picard-Lefschetz methodology Marolf:2022ybi ; Chen:2025leq (see Witten:2010cx ; Feldbrugge:2017kzv ; Narain:2021bff ; Shoji:2025riv ; Tanizaki:2014xba ; Witten:2010zr for discussion on Picard-Lefschetz technique).
@@
The essence of this procedure is that we do not deal with the conformal factor problem of the Euclidean path-integral, instead we use the Lorentzian path-integral to define a Euclidean one via an integral transform, which, for the case of black hole geometries, reduces to the known partition function.
Once we have comfortably defined the Euclidean partition function $Z(\beta)$ for the case of black hole geometries, we proceed to study it in detail for the case of AdS-Schwarzschild and charged black holes in arbitrary dimensions. Our starting point in this study is eq. (7) and eq. (8). These are either one or two-dimensional integral which can be evaluated by analysing them in the complex plane. As this targets geometries which correspond to black hole solutions, it is worth exploring what are the possible saddles of the system and what do they correspond to. Furthermore, it is crucial to investigate which saddles give dominant contribution to the partition function for various values of parameter $\beta,Q,\Phi$ and whether they are compatible with the known results of AdS/CFT for various dimensions.
It may happen that certain saddles, which are dominant in the partition function for various ranges of parameters, are incompatible with the known wisdom of AdS/CFT. In such cases, it is worth asking whether such saddles become “relevant” in the partition function when the integrations in eqs. (7) and (8) are performed in the complex plane using Picard-Lefschetz methods. An interesting situation seen in the recent work Mahajan:2025bzo ; Singhi:2025rfy is that there are saddles that are at odds with the usual expectations from AdS/CFT, do not contribute to the partition function. They are “not” relevant according to the PL method. This paper investigates such situations in more detail for uncharged and charged black holes in arbitrary $d+1$-dimensions.
The outline of the paper is as follows: In Sec. II, we revisit the puzzle and findings reported in Mahajan:2025bzo for AdS Schwarzschild black hole by generalizing to arbitrary dimensions. Here, we discuss the implementation of Picard Lefschetz and homology averaging of thimbles, along with thermodynamic stability from the viewpoint of PL analysis. In Sec. III, we extend this study to the charged AdS Schwarzschild black holes both in grand canonical ensemble (fixed potential) and the canonical ensemble (fixed charge). Sec. IV is devoted to the analysis of weaker verison of the KSW criterion and its compatibility with the puzzle raised in the earlier sections. Finally, summary of our findings and conclusions are presented in section V.
II AdS Schwarzschild black hole in $d+1$ dimension
@@
This is solution to equation of motion following from Einstein-Hilbert action with appropriate surface terms and counterterms to cancel the large volume divergence Skenderis:2002wp . For $\mu=0$, it describes a thermal AdS solution without any horizon. However, when $\mu>0$, we get a horizon at $r_{+}$, which is the largest positive real root of $f(r_{+})=0$ and $r_{+}\leq r<\infty$. Here, $r\geq r_{+}$ ensures the existence of a timelike killing vector everywhere outside the horizon, and regularity at the horizon enforces $\beta$ to be $4\pi/f^{\prime}(r_{+})$. The entropy $S$ and energy $E$ as function of $r_{+}$ are given by
$S(r_{+})=\frac{\omega r_{+}^{d-1}}{4G},\hskip 11.38109ptE(r_{+})=\frac{\omega(d-1)}{16\pi G}\left(r_{+}^{d-2}+\frac{r_{+}^{d}}{l^{2}}\right)\,$, (10)
@@
$\beta_{\rm max}=\frac{2l\pi}{\sqrt{d(d-2)}}\,.$ (14)
@@
$\begin{split}K(r)=\frac{d(d-1)^{2}(d-2)\mu^{2}}{r^{2d}}+\frac{2d(d+1)}{l^{4}}\,,\quad{\rm where}\quad\mu^{2}=r_{+}^{2(d-2)}\biggl(1+\frac{r_{+}^{2}}{l^{2}}\biggr)^{2}\,.\end{split}$ (17)
From eq (17), it is evident that at $r=0$, $K(r=0)$ blows up in $d\geq 3$, signifying the presence of curvature singularity irrespective of whether $r_{+}$ is real (for $\beta<\beta_{\rm max}$) or complex (which happens for $\beta>\beta_{\rm max}$). The real saddles describe the Euclidean cigar geometries, while in the Lorentzian continuation, they correspond to a black-hole with a singularity hidden behind the horizon. On the other hand, for complex saddles things are different. Their Lorentzian continuation correspond to naked singularities. Such naked singularities are also solutions of Einstein’s equations.
In the following subsections, we will study the partition function for various regimes of $\beta$, analysing the saddle configurations, their dominance and relevance, and compute the partition function using the methods of Picard-Lefschetz. To do so, we would need to study the argument of the exponent in eq. (11) at the various saddles. $\mathcal{I}(r_{+})$ at various saddles is given by
$\begin{split}&\mathcal{I}(r_{+}=0)=0\,,\\ &\mathcal{I}(r_{+}=r_{+}^{\pm})=\frac{\omega l^{d-1}}{4G}\frac{(d-2)^{(d-3)/2}}{d^{(d+1)/2}}\exp\{{\pm(d-1)\theta}\}\bigl[-1\pm(d-1)\tanh\theta\bigr]\,.\end{split}$ (18)
@@ II.1 $Z(\beta)$ for $\beta>\beta_{\rm max}$
In this regime, we have one real saddle and two complex saddle configurations. The real saddle corresponds to thermal AdS geometry, while the complex-conjugate saddle correspond to naked singular geometries. As noticed in Mahajan:2025bzo , these complex saddles give dominant contribution in the partition function for certain range of values of $\beta$, as can be seen by computing $\mathcal{I}(r_{+})$. Naively, it will be expected that being dominant saddles, they are the ones whose contribution will be most important in the semi-classical limit. However, the situation changes if such saddles, despite being dominant becomes irrelevant in the computation of the partition function using Picard-Lefschetz methods. This situation was analysed for the case of $d=4$ in Mahajan:2025bzo . In our paper we examine the situation in arbitrary dimensions.
In order to compute the partition function for $\beta>\beta_{\rm max}$, we start by analysing ${\rm Re}(\mathcal{I}(r_{+}))$ at various saddles. It will indicate the saddles which will become dominant as $\beta$ varies. It is seen from eq. (18) that for arbitrary $d$, $\mathcal{I}(r_{+})$ for thermal AdS is always zero. For $\beta>\beta_{\rm max}$, the other two saddles $r_{+}^{\pm}$ are complex conjugate and correspond to naked singularities. For them $\mathcal{I}(r_{+})$ is complex as seen from eq. (18) . To analyse the dominance, one needs to extract the real part of ${\rm Re}[\mathcal{I}(r_{+})]$ for each of these complex saddles. If the ${\rm Re}[\mathcal{I}(r_{+})]>0$ for certain values of $\beta$, then these naked singularities become dominant over the thermal-AdS saddle, which will be in conflict with the results of AdS/CFT. For $\beta>\beta_{\rm max}$, the ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]$ is given by
@@
It is easy to see from here that for certain values of $\beta$, say for example large-$\beta$ ($\beta\to\infty$), $\cos((d-1)\theta)=0$ or $\pm 1$, while $\sin((d-1)\theta)=\pm 1$ or 0 (depending on the dimension $d$), respectively. In the case when it is 1-1 (for example, $d=4$), then naked singularities are dominant in the partition function! More generically, for large-$\beta$, when $d=4n,n\in(1,2,3,..)$, ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]\sim+(\beta/\beta_{\rm max})>0$ thereby claiming dominance over thermal-AdS. This was already noticed in Mahajan:2025bzo for $AdS_{5}(n=1)$. When d=4n1d=4n-1, for large-$\beta$, ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]<0$ (independent of $\beta$) and thermal-AdS remains a dominant saddle. However, for $d=4n+1$, ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]>0$ and is independent of $\beta$, thereby again dominating over the thermal-AdS saddle. For $d=4n+2$, ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]\sim-(\beta/\beta_{\rm max})$ for large-$\beta$ thereby making thermal-AdS dominant saddle. We summarize this in the table 1.
@@
Table 1: Dominant saddles of the partition function in the large $\beta(\gg\beta_{\rm max})$ limit for various dimensions d with $n\in(1,2,3,\cdots)$.
On doing a numerical plot of ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]$ as a function of arbitrary $\beta>\beta_{\rm max}$, we notice from fig. (1) the values of $\beta$ for which naked singularities give dominant contribution in the partition function.
Figure 1: Real part of $\mathcal{I}$ (see eq. (19)) as a function of $\beta$ ($\beta_{\rm max}$) at various saddles for $d=4$ and $d=7$. ${\rm Re}[\mathcal{I}(r_{+}^{\pm})]>0$ for complex saddles (naked-singular geometries) imply that for these values of $\beta$ naked singularities are dominant over the Thermal-AdS saddle.
From this discussion, it is clear that for certain values of $\beta$ for the case of $\beta>\beta_{\rm max}$, naked singular geometries, which are complex saddles, are dominant. Naively, one would expect that they would control the nature of the partition function in the semiclassical limit, as they are most dominant, though they correspond to complex metrics ($r_{+}$ being complex implies $\mu$ is complex). Occurrences of complex metrics are not new in the literature. They have shown up in studies of quantum cosmology, black holes Feldbrugge:2017kzv ; Maldacena:2024uhs ; Ivo:2024ill ; Ailiga:2024mmt ; Ailiga:2023wzl ; DiTucci:2020weq ; Lehners:2021mah ; Jonas:2022uqb ; Nishimura:2024one ; Hertog:2015nia ; Cabo-Bizet:2020ewf and real-time holography as well Skenderis:2008dh ; Glorioso:2018mmw ; Jana:2020vyx , where they have lead to physically meaningful results. However, in our case if they are to be considered, then our findings will be in tension with the understanding of black hole thermodynamics. Furthermore, if these additional saddles are considered, new phase transitions will occur among these complex saddles. This will also be in conflict with AdS-CFT, where it is well understood that thermal AdS reproduces results consistent with the dual CFT in the low temperature regime Maldacena:1997re ; Witten:1998zw ; Witten:1998qj ; Aharony:2003sx .
This is the puzzle first raised in the note Mahajan:2025bzo , where naive computation of the partition function based on saddle dominance lead to tension with AdS/CFT. However, if the partition function is computed carefully, making use of Picard-Lefschetz methods and tools from Morse theory, then the puzzle is resolved on its own, as noticed in Mahajan:2025bzo . In this paper, we study this situation for generic dimensions and analytically show that complex saddles corresponding to naked-singular geometries become irrelevant in the path-integral and hence don’t contribute.